DIRC Operations Weekly Report
Period: Wednesday 2007/4/10 - Wednesday 2007/4/17
Oncall: Alejandro Perez
Fast Monitoring: Alejandro Perez, Bradley Wogsland
PEP peak lumi: ~1. x 1034
(previous week: ~8.3 x 1033)
BaBar integrated lumi: 3.28 fb-1
(previous week: 0.80 fb-1)
Run 6 PEP delivery: 3.43 fb-1
(previous week: 0.92 fb-1)
Run 6 total delivery: ~28.42 fb-1
Run 6 total BaBar recorded: 27.25 fb-1
including 0 fb-1 off-peak data
PEP activities this week:
2007/04/13: Problems during the swing shift, PEP was trying to put HER RF 12-6 online, and the same problem for LER RF 4-2 during owl.
2007/04/16: Problem during night in the BSY region.
2007/04/17: In the morning, PEP needed access to region 4 to fix magnet problems.
BaBar activities/issues this week:
- Since 2007/04/13 power Chiller SIAM glitches. Duration from 0.25sec to 2.0sec occuring at random.
logbook for details
- ROD next thursday. Ray is going to scope the power DIRC chiller and SIAM to investigate the power chiller SIAM glitches since 2007/04/13.
- Replace two DFBs, one in sector 0 and one in sector 1 with a group of 8 innefficient or
flaky channels (require bridge).
- Replace new DFB in sector 8 with 18 (TDC chip 0) often dead channels and figure
out what is killing DFBs in that slot.
- Found a mini X-tree in sector 7 HV group 55 than only shows up during taking cosmics.
Jose, Alejandro and Bradley rotating backup & on-call commissioning